My homework is asking to find the surface area of this prism, rounded to the nearest whole unit. Here's the problem:
1. A triangular prism with base perimeter 24 cm, base area 24 cm2, and height 15 cm.

Answer by toidayma(44)

The surface area of a triangular prism equal = 2*base area (there are 2 bases) + surround area (area of three rectangles)
Bases' area = 2*24 = 48 (cm2)
About the area of the three surround rectangles, Imagine that these three rectangles have a size with same length (the height of the prism). The left sides are three sides of the triangular base. You don't know the length of each side, but you do know the sum of all three sides, which is the perimeter of the base.
Let a,b,c are length of three sides of the triangular base.
The area of the three surround rectangles are a*h, b*h, and c*h. Therefore, the area of the surround surface (three rectangles) is h(a+b+c) = h*(base perimeter)= 15*24 = 360 (cm2)
It is easy from this point. The surface area = 48 + 360 = 408 (cm2).